OpenStudy (anonymous):

(10)/(3+x)+(x+1)/(3+x)

OpenStudy (anonymous):

(x+11)/(3+11)

OpenStudy (anonymous):

but how you did it?

OpenStudy (anonymous):

Can you explain to me how you did that?

OpenStudy (mimi_x3):

\[\large \frac{10}{3+x} +\frac{x+1}{3+x} = \frac{10+x+1}{3+x} = \frac{11+x}{3+x} \]

OpenStudy (anonymous):

I see,your answer is wrong,you didn't follow the rules on how to add and subtract rational algebraic expression.

OpenStudy (mimi_x3):

What..? My answer is not wrong. There's no substraction involved in this question. http://www.wolframalpha.com/input/?i=%2810%29%2F%283%2Bx%29%2B%28x%2B1%29%2F%283%2Bx%29